1. now we look at the role taxes play in determining equilibrium income. Suppose we have an economy described by the following functions:


C = 500 + 0.8YD


I  = 700


G = 300


TR = 100

TA =  400

NX = -100


a. Calculate the equilibrium level of income and the multiplier in this model.

b. Calculate also the budget surplus, BS.

c. Suppose that TA increases to 500. What is the new equilibrium income? 




1
Expert's answer
2022-03-14T10:30:34-0400

a. Calculate the equilibrium level of income and the multiplier in this model.

At equilibrium, YD=I+C

YD=700+500+0.8YD

YD-0.8YD=1200

"\cancel{0.2}YD=\frac{1200}{0.2}=6,000"

Equilibrium income=

Income multiplier=0.8


b. Calculate also the budget surplus, BS.

S=YD-C-G-T-Nx=6,000-980-700-300+400=4,420

c. Suppose that TA increases to 500. What is the new equilibrium income? 

New equilibrium=6000-100=5,900